OpenStudy (anonymous):

Evaluate when t=1.5 [(t+2)squared-3t]/5 the answer is (t squared +t+4)/5 right?

OpenStudy (precal):

are you just plugging in t=1.5? This sounds like particle motion (calculus and physics).

OpenStudy (anonymous):

I just plugged it into the simplified form and got 1.55. Is that right?

OpenStudy (precal):

Is this calculus? Particle motion is not that easy.

OpenStudy (anonymous):

no i think it's algebra

OpenStudy (precal):

Ok then let's do it using Algebra only \[\left[ \left( t+2 \right)^2-3t \right]/5\] if t=1.5 then \[\left[ \left( 1.5+2 \right)^2-3*1.5 \right]/5\] \[\left[ \left( 3.5 \right)^2-4.5 \right]/5\] \[\left( 12.25-4.5 \right)/5\] \[\left( 7.75 \right)/5\] 1.55

OpenStudy (anonymous):

grazi, wanna help me with another one, simpler ?

OpenStudy (precal):

Go for it! If I can help you, I will.

OpenStudy (anonymous):

i'm posting it in the chat

OpenStudy (precal):

by the way you answer is correct if you did not have t=1.5

OpenStudy (anonymous):

huh???

OpenStudy (precal):

\[\left( (t+2)^2-3t \right)/5=\left( t^2+t+4 \right)/5\]